Determine the amount of money that must be invested now​ (time 0) at 7​% nominal​ interest, compounded​ monthly, to provide an annuity of ​$5,000 per year for 12 ​years, starting eight years from now. The interest rate remains constant over this entire period of time.

The amount of money that must be invested now is ​$
Answer #1

First we need to calculate the effective rate of interest.

Effective rate of interest=i=(1+7%/12)^12-1=7.2290%

Amount of investment should be equal to PV of future cash inflows i.e.

Amount of investment now=5000*(P/A,0.072290,12)*(P/F,0.072290,8)

Let us calculate the interest factors

(P/F,0.072290,8)=1/(1+0.072290)^7=0.61349949

Amount of investment now=5000*7.84664826*0.61349949=$24069.57